Morrisey & Brown, Ltd., of Sydney is a merchandising company that is the sole distributor of a product that is Increasing in popularity among Australlan consumers. The companys Income statements for the three most recent months follow: Norrisey & B Income Statements For the Three Months Ended Septenber 30 gus 7,000 7.500 Salem in unita Sales Cost of goods sold Gross margin Selling and administrative expenses i 6,500 s650,00070000 75,000 450,000 390,000 420,000 260,000280,000300,000 Advertising expense Shipping expense Salaries and commiasions Insurance expense Depreciation expense 25,800 69,000 118,000 9,750 16,400 238.950 $ 21,050 25,800 73,000 125,000 9,750 16,400 249.950 25,800 77,000 132,000 9,750 16,400- Total selling and administrative expenses Net operating income 260,950 39,050 30,050 Required: 1. By analyzing the data from the companys income statements, classify each of its expenses (including cost of goods sold) as either variable, fixed, or mixed 2. Using the high-low method, separate each mixed expense into variable and fixed elements. Express the variable and fixed portions of each mixed expense in the form Y=a + bX. 3. Redo the companys income statement at the 7,500-unit level of activity using the contribution format.
Answer #1

1.

Cost item Nature
Cost of goods sold Variable
Advertising expenses Fixed
Shipping expense Mixed
Salaries and commissions Mixed
Insurance expense Fixed
Depreciation expense Fixed

2.

Variable cost per unit = (High cost - Low cost)/Difference in units

Hence, variable shipping expense = (77,000 - 69,000)/1,000

= $8

Variable shipping expense at 7,500 units = 7,500 x 8

= $60,000

Total shipping expense at 7,500 units = $77,000

Hence, fixed shipping expense = 77,000 - 60,000

= $17,000

Cost function for shipping expense (Y) = 17,000 + 8X

Variable salaries and commission expense = (132,000 - 118,000)/1,000

= $14

Variable salaries and commission expense at 7,500 units = 7,500 x 14

= $105,000

Total salaries and commission expense at 7,500 units = $132,000

Hence, fixed salaries expense = 132,000 - 105,000

= $27,000

Cost function for salaries and commission expense (Y) = 27,000 + 14X

3.

Income statement

Sales 750,000
Variable cost:
Cost of goods sold 450,000
Shipping expense 60,000
Salaries and commission expense 105,000
Total variable cost - 615,000
Contribution margin 135,000
Fixed costs:
Shipping expense 17,000
Salaries and commission expense 27,000
Advertising expense 25,800
Insurance expense 9,750
Depreciation expense 16,400
Total fixed costs - 95,950
Net operating income 39,050

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.